OpenStudy (ammarah):

Write a rule for the nth term of the geometric sequence given the info. a1=4, a2=12

OpenStudy (aravindg):

We need more info. It can be either an AP or GP.

OpenStudy (ammarah):

whats ap of gp mean?

OpenStudy (aravindg):

AP mean arithmetic progression. GP mean Geometric progression.

OpenStudy (ammarah):

i know the question states its a geometric progression....

OpenStudy (primeralph):

a2/a1 = 3, so 4*3^(n-1).

OpenStudy (ammarah):

? idk what ur saying...

OpenStudy (ammarah):

a2 divided by a1????

OpenStudy (ammarah):

pleasde help im dying.

OpenStudy (ayubie):

since its geometric, try to figure out how to get from 4 to 12 geometrically (multiply by something)

OpenStudy (ammarah):

so times 3

OpenStudy (ayubie):

yeah. using the formula An = A1 * r ^ (n-1). we just found out that r is 3, and you already know A1 is 4. The answer given by primeralph was correct, when you plug in all of the values you get that An = 4*3^(n-1)

OpenStudy (ammarah):

oohk so thats the answer?

OpenStudy (ammarah):

How would u do this: Write the series -5+5+15+25+35 using summation notation?

OpenStudy (ammarah):



OpenStudy (ammarah):

@ayubie

OpenStudy (ayubie):

OHH. okay haha give me just a minute.

OpenStudy (ammarah):